# Differentially-Private Distributed Model Predictive Control of Linear Discrete-Time Systems with Global Constraints

### v1

@@ 5 Numerical Simulations
For all subsystems, the local state and input constraint sets are selected as $\mathcal{X}_{i}=\{x_{i}:-1\leq x_{i}\leq 1\}$ and $\mathcal{U}_{i}=\{u_{i}:-0.3\leq u_{i}\leq 0.3\}$, respectively. The global constraint is $-0.65\leq\sum_{i=1}^{4}u_{i}\leq 0.65$. The weight matrices $Q_{i}$ and $R_{i}$ are set as $Q_{i}=I$ and $R_{i}=0.1$, respectively. The length of the prediction horizon is chosen as $N=5$. In Algorithm 3, we inject Laplace noise with parameter $\nu^{k}=0.1+0.001k^{0.1}$. The weakening factor sequence and step-size sequence is set as $\chi^{k}=\frac{2}{1+0.01k^{0.9}}$ and $\gamma^{k}=\frac{5}{1+0.1k}$, respectively. In the simulation, we run Algorithm 4 for 20 times and calculate the average and the variance of the state and input trajectories. For comparison, we also run Algorithm 2 under the same noise level.
@@
The simulation results are illustrated in Figures 2 and 3. Figure 2 depicts the evolution of the system state. It can be seen that the variance of the system state trajectories under Algorithm 2 is much larger than those under Algorithm 4. This discrepancy arises from the direct integration of persistent noise into Algorithm 1, which sacrifices optimization accuracy and subsequently degrades the control performance of Algorithm 2. In our developed approach, the weakening factor $\chi^{k}$ is tailored to alleviate the influence of DP noise, ensuring accurate convergence of the system state. In addition, Figure 3 presents the evolution of the global constraint. It can be found that there exist constraint violations in Algorithm 2. However, owing to the implementation scheme developed in Section 4, our approach can guarantee the satisfaction of the global constraint.
@@
6 Conclusion
This paper developed a differentially private DMPC strategy for linear discrete-time systems with coupled global constraints. We showed that the DMPC method relying on the conventional distributed dual-gradient algorithm is susceptible to eavesdropping attacks. To address this issue, we incorporated a DP noise injection mechanism into the distributed dual-gradient algorithm, enabling privacy preservation while maintaining accurate optimization convergence. Furthermore, a practical implementation approach for DMPC was proposed, which guarantees the feasibility and stability of the closed-loop system. Simulation results validated the effectiveness of the proposed DMPC strategy.

### v2

@@ V Numerical Simulations
For all subsystems, the local state and input constraint sets are selected as $\mathcal{X}_{i}=\{x_{i}:-1\leq x_{i}\leq 1\}$ and $\mathcal{U}_{i}=\{u_{i}:-0.3\leq u_{i}\leq 0.3\}$, respectively. The global constraint is $-0.65\leq\sum_{i=1}^{4}u_{i}\leq 0.65$. The weight matrices $Q_{i}$ and $R_{i}$ are set as $Q_{i}=I$ and $R_{i}=0.1$, respectively. The length of the prediction horizon is chosen as $N=5$. In Algorithm 3, we inject Laplace noise with parameter $\nu^{k}=0.1+0.001k^{0.1}$. The weakening factor sequence and step-size sequence is set as $\chi^{k}=\frac{2}{1+0.01k^{0.9}}$ and $\gamma^{k}=\frac{5}{1+0.1k}$, respectively. In the simulation, Algorithm 4 is executed 20 times, and the mean and the variance of the state and input trajectories are computed. For comparison, we also run Algorithm 2 under the same noise level.
@@
The simulation results are illustrated in Figures 2 and 3. Figure 2 depicts the evolution of the system state. It can be seen that the variance of the system state trajectories under Algorithm 2 is much larger than those under Algorithm 4. This discrepancy arises from the direct integration of persistent noise into Algorithm 1, which sacrifices optimization accuracy and subsequently degrades the control performance of Algorithm 2. In our developed approach, the weakening factor $\chi^{k}$ is tailored to alleviate the influence of DP noise, ensuring accurate convergence of the system state. In addition, Figure 3 presents the evolution of the global constraint. It can be found that there exist constraint violations in Algorithm 2. However, owing to the implementation scheme developed in Section IV, our approach can guarantee the satisfaction of the global constraint.
@@
VI Conclusion
This paper developed a differentially private DMPC strategy for linear discrete-time systems with coupled global constraints. We showed that the DMPC method relying on the conventional distributed dual-gradient algorithm is susceptible to eavesdropping attacks. To address this issue, we incorporated a DP noise injection mechanism into the distributed dual-gradient algorithm, enabling privacy preservation while maintaining accurate optimization convergence. Furthermore, a practical implementation approach for DMPC was proposed, which guarantees the feasibility and stability of the closed-loop system. Simulation results validated the effectiveness of the developed privacy-preserving DMPC strategy. Future work will extend the differentially private framework for systems with uncertainties (e.g., robust and stochastic DMPCs).